A software designer is mapping the streets for a new racing game. All of the streets are depicted as either perpendicular or parallel lines. The equation of the lane passing through A and B is -7x + 3y = -21.5. What is the equation of the central street PQ?


1
Expert's answer
2021-07-14T18:16:12-0400

The equation of the lane passing through "A" and "B," is


"-7x+3y=-21.5"

Solve for "y"


"y=\dfrac{7}{3}x-\dfrac{43}{6}"

Hence

"slope_1=\dfrac{7}{3}"

Since the central street "PQ" is perpendicular to the lane passing through "A" and "B," then


"slope_2=\dfrac{-1}{slope_1}=-\dfrac{3}{7}"

Point "P(7, 6)"


"6=-\dfrac{3}{7}(7)+b=>b=9"

The equation of the central street "PQ" is


"y=-\dfrac{3}{7}x+9"